# OmnimatteZero: Training-free Real-time Omnimatte with Pre-trained Video Diffusion Models

### v1

@@ 4.1 Removing associated object effects
We are interested in masking the object with its associated effects. Recently, [15] showed that pretrained video diffusion models can associate objects with their effects by analyzing self-attention maps between query and key tokens related to the object of interest. They leveraged this insight to train diffusion models for object removal along with its associated footprint. In contrast to [15], we propose to directly derive the masks from the attention maps, allowing a training-free object removal approach. Specifically, we apply a single noising and denoising step on the video latent and compute the attention weights between query and key tokens associated with the object. More precisely, for each latent frame $i$ at diffusion layer $l$, we calculate the attention map as follows:
@@
$A_{i}^{l}=\text{softmax}\left(\frac{Q_{i}^{l}(M_{obj}\odot K_{i}^{l})^{T}}{\sqrt{d}}\right)\in{\mathbb{R}}^{w\times h}$ (2)
@@
Figure 2: Self-attention maps from (a) LTX Video diffusion model and (b) Image based Stable Diffusion. The spatio-temporal video latent ”attends to object associated effects” (e.g., shadow, reflection) where, image models struggles to capture these associations.
We then compute a soft-mask per-frame, $\mathcal{M}_{i}\in[0,1]^{w\times h\times c}$, which is aggregated across all diffusion layers:
$\mathcal{M}_{i}=\frac{1}{L}\sum_{l=1}^{L}A_{i}^{l}$ (3)
where $L$ is the number of layers in the video diffusion model. Then we extend $\mathcal{M}_{i}$ to the whole video by concatenating all frame masks channel-wise, deriving a soft-mask latent code $\mathcal{M}_{obj}=\text{concat}\{\mathcal{M}_{1},...\mathcal{M}_{f}\}\in{\mathbb{R}}^{w\times h\times f}$. To obtain a binary mask we perform Otsu thresholding [23] for each latent frame, $\hat{M}_{obj}=\mathbb{I}(\mathcal{M}_{obj})$. This new mask replaces the mask $M_{obj}$ provided as input for the two inpainting methods and the soft mask $\mathcal{M}_{obj}$ is used for the latent blending in Equation 1 a process we now refer to as attention-based latent blending. We found that soft blending the two latents leads to improved results, as it allows the diffusion model to leverage spatial and temporal information from both latents more effectively during the final denoising stage. Fig. 2(a) visualizes these self-attention maps (obtained from LTX-Video [9]) to two videos, demonstrating the success of the attention map, in capturing the object and its associated traces, e.g. the cat and its reflection and the bicycle rider with its shadow.
Interestingly, to the best of our knowledge, this approach has not been explored for masking object effects (e.g. shadows) in images. Unlike video diffusion models, image models do not capture object effects from still images [32]. Figure 2(b) illustrates this by showing self-attention maps extracted using StableDiffusion [26], a text-to-image diffusion model, which demonstrates that the object does not attend to its associated effects. We believe that this aligns with the principle of common fate in Gestalt psychology [13], which suggests that elements moving together are perceived as part of the same object. Video diffusion models seem to implicitly leverage this principle by grouping objects with their effects through motion.
Figure 3: (a) Foreground Extraction: The target object is extracted by latent code arithmetic, subtracting the background video encoding from the object+background latent (Latent Diff). This initially results in distortions, which are later corrected by replacing pixel values using the original video and a user-provided mask (Latent Diff + Refinement). (b) Layer Composition: The extracted object layer is added to a new background latent (Latent Addition). To improve blending, a few steps of SDEdit are applied, yielding a more natural integration of the object into the new scene (Latent Addition + SDEdit).
@@
ObjectDrop [32] 28.05 0.124 34.22 0.083 31.14 0.104 - -
Video RePaint [LTXVideo]* [21] 20.13 0.252 21.15 0.289 20.64 0.271 0 0.04
Video RePaint [Wan2.1]* [21] 21.44 0.244 24.16 0.261 22.8 0.253 0 3.2
Lumiere inpainting [2] 26.62 0.148 31.46 0.157 29.04 0.153 - 9
Propainter [36] 27.44 0.114 34.67 0.056 31.06 0.085 - 0.083
DiffuEraser [36] 29.51 0.105 35.19 0.048 32.35 0.077 - 0.8
Ominmatte [19] 21.76 0.239 26.81 0.207 24.29 0.223 3 2.5
D2NeRF [33] - - 34.99 0.113 - - 3 2.2
LNA [12] 23.10 0.129 - - - - 8.5 0.4
OmnimatteRF [17] 33.86 0.017 40.91 0.028 37.38 0.023 6 3.5
Generative Omnimatte [15] 32.69 0.030 44.07 0.010 38.38 0.020 - 9
OmnimatteZero [LTXVideo] (Ours) 34.11 0.015 44.07 0.010 39.09 0.012 0 0.04
OmnimatteZero [Wan2.1] (Ours) 33.89 0.019 44.25 0.008 39.02 0.013 0 3.2
Table 1: Quantitative comparison: Background Reconstruction. OmnimatteZero outperforms all omnimatte and video inpainting methods, achieving the best PSNR and LPIPS without training or per-video optimization. It also runs significantly faster, with OmnimatteZero [LTXVideo] at 0.04s per frame. ”-” denotes missing values due to unreported data or unavailable public code.
4.2 Foreground extraction
We can now use our object removal approach to extract object layers along with their associated effects. To isolate a specific object, we apply our method twice: first, removing all objects from the video, leaving only the background, denoted as $V_{bg}\in\mathbb{R}^{F\times W\times H\times 3}$ with its corresponding latent $Z_{bg}$; second, removing all objects except the target object, resulting in a video of the object with the background, denoted as $V_{obj+bg}\in\mathbb{R}^{F\times W\times H\times 3}$ with its corresponding latent $Z_{obj+bg}$. We can now derive the video of the object isolated from the background by simply subtracting the two latents $Z_{obj}=Z_{obj+bg}-Z_{bg}$. Applying thresholding on $Z_{obj}$ results in a latent that is decoded to a video $V_{obj}$ containing only the object and its associated effects (see Figure 3a Latent Diff). While the extracted effects appear convincing, the object itself often suffers from distortions due to latent values falling outside the diffusion model’s learned distribution. To address this issue, we make use of the information available in the pixel-space. We refine the object’s appearance by replacing its values in the pixel-space with those from the original video, based on the user provided mask. Specifically,
$V_{obj}=M^{p}_{obj}\odot V_{obj+bg}+\left(1-M^{p}_{obj}\right)\odot V_{obj}.$ (4)
@@ 4.3 Layer composition
5 Results
Following [15], we evaluate OmnimatteZero on three applications: (1) Background Reconstruction, where the foreground is removed to recover a clean background; (2) Foreground Extraction, where objects are extracted, together with their associated effects (e.g. shadows and reflections); and (3) Layer Composition, where extracted elements are reinserted into new backgrounds while preserving visual consistency.
Implementation Details:
We apply OmnimatteZero using two video diffusion models: LTXVideo [9] v0.9.1 and Wan2.1 [29], both running with $T_{\text{steps}}=50$ denoising steps. For all experiments, we set $k=45$, as we found that just five steps are sufficient for blending. Input videos were resized to $(768,512)$ before processing. The guidance scale was set to 0 (i.e. no prompt), and all result videos were generated using the same random seed.
Quantitative and qualitative results of baselines were obtained from the respective authors. However, since only OmnimatteRF’s code is publicly available, we compare against all methods when possible but limit comparisons to OmnimatteRF for new videos. All qualitative results in this paper are based on LTXVideo, as it shows minimal visual differences from Wan2.1. Additionally, due to space constraints, each figure displays a single frame per video for qualitative results. Full videos are available in the supplementary material.
Figure 4: Qualitative Results: Object removal and background reconstruction. The first row shows input video frames with object masks, while the second row presents the reconstructed backgrounds. Our approach effectively removes objects while preserving fine details, reflections, and textures, demonstrating robustness across diverse scenes. Notice the removal of the cat’s reflection in the mirror and water, the shadow of the dog and bicycle (with the rider), and the bending of the trampoline when removing the jumpers.
Figure 5: Qualitative Comparison: Object removal and background reconstruction. Our approach achieves cleaner background reconstructions with fewer artifacts while Generative Omnimatte [15] leaves some residuals, DiffuEraser [16] and ProPainter [36] struggle with noticeable traces (highlighted in red). The last two rows show that OmnimatteZero successfully removes effects where others fail.
5.1 Background layer reconstruction
@@
We conducted a comparative evaluation of our approach with several SoTA methods. These methods fall into three categories: (A) Omnimatte methods that are trained to decompose a given video into layers: Omnimatte [19], D2NeRF [33], LNA [12], OmnimatteRF [17] and Generative Omnimatte [15]. (B) Video inpainting methods that are trained to remove objects from a video: RePaint [21] adapted for video, Lumiere inpainting [2], Propainter [36] and DiffuEraser [16]. Finally, (C) An image inpainting method that is applied for each video frame independently: ObjectDrop [32].
Datasets and Metrics:
@@
We evaluate our method on two standard Omnimatte datasets: Movies [17] and Kubric [33]. These datasets provide object masks for each frame and ground truth background layers for evaluation. All methods are assessed using PSNR and LPIPS metrics to measure the accuracy of background reconstruction, along with comparisons of training and runtime efficiency on a single A100 GPU.
Quantitative results:
@@
Table 1 presents a comparison of OmnimatteZero with the SoTA Omnimatte and inpainting methods on the Movies and Kubric benchmarks. It shows that OmnimatteZero outperforms existing supervised and self-supervised methods designed for object inpainting or omnimatte. It achieves the highest PSNR and lowest LPIPS on both the Movie and Kubric datasets, demonstrating superior background reconstruction. Specifically, OmnimatteZero [LTXVideo] achieves a PSNR of 34.11 (Movie) and 44.07 (Kubric), surpassing Generative Omnimatte [15], all while requiring no training or per-video optimization. Notably, this improvement is not due to a stronger video generation model, as both OmnimatteZero and Video RePaint [21] use the same generator, yet Video RePaint records the lowest PSNR and highest LPIPS across all benchmarks.
@@
Our method is also significantly faster. OmnimatteZero [LTXVideo] runs at just 0.04s per frame (or 24 frames per second). These results establish OmnimatteZero as the first training-free, real-time video matting method, offering both superior performance and practical efficiency.
@@
Figure 6: Qualitative Comparison: Foreground Extraction. Foreground extraction comparison between OmnimatteZero and OmnimatteRF [17]. Our method accurately captures both the object and its associated effects, such as shadows and reflections, in contrast to OmnimatteRF, often missing or distorting shadows (row 2) and reflections (row 3).
Qualitative results:
@@
Figure 4 presents qualitative results of our object removal method across various scenes. The first row shows input video frames with masked objects, while the second row displays the reconstructed backgrounds. The videos include (1) a cat running toward a mirror, (2) a static cat looking left and right, (3) a dog running toward a man, (4) a cyclist passing, and (5) multiple people jumping on a trampoline. Our method effectively removes objects like people, animals, and bicycles while preserving fine details and textures. Notably, in the first two columns, OmnimatteZero eliminates the cat without leaving mirror or water reflections. The last column further demonstrates its ability to remove multiple moving objects while maintaining scene integrity, even correcting the trampoline’s bending effect after removing the jumper.
Figure 5 presents a qualitative comparison of OmnimatteZero with SoTA video matting and inpainting methods. Our approach consistently produces cleaner background reconstructions with fewer artifacts.
The videos in the figure include (1) penguins walking on the shore, (2) a car passing on a wet road, (3) a girl playing with Lego bricks, and (4) people jumping on a trampoline. In the first two rows, both OmnimatteZero and Generative Omnimatte [15] successfully remove the penguins and car, naturally filling the background without noticeable artifacts. In contrast, DiffuEraser [16] and ProPainter [36] struggle, leaving visible artifacts (highlighted in red). The third and fourth rows emphasize OmnimatteZero’s advantage, effectively removing effects like shadows, while other methods fail, leaving shadows or generating blur artifacts.
@@
Figure 7: Qualitative Comparison: Layer Composition. The extracted foreground objects, along with their shadows and reflections, are seamlessly integrated into diverse backgrounds, demonstrating the effectiveness of our approach in preserving visual coherence and realism across different scenes.
5.2 Foreground extraction
@@
We aim to extract the foreground object along with its associated effects, such as shadows and reflections. Figure 6 compares OmnimatteZero for foreground extraction with OmnimatteRF [17]. Our training-free method accurately isolates objects while preserving fine details of their interactions with the scene, such as reflections and shadows. In contrast, OmnimatteRF struggles to fully capture the associated effects. Notably, in the second and third rows, our method correctly extracts both the shadow of the cyclist and the reflection of the cat, while OmnimatteRF either distorts or omits these elements. These results demonstrate that OmnimatteZero provides superior, training-free foreground extraction, effectively handling complex visual effects where existing methods fall short.
5.3 Layer composition

### v2

@@ 5.2. Removing associated object effects
We are interested in masking the object with its associated effects. Recently, (Lee et al., 2024) showed that pretrained video diffusion models can associate objects with their effects by analyzing self-attention maps between query and key tokens related to the object of interest. They leveraged this insight to train diffusion models for object removal along with its associated footprint. In contrast to (Lee et al., 2024), we propose to directly derive the masks from the attention maps, allowing a training-free object removal approach. Specifically, we apply a single noising and denoising step on the video latent and compute the attention weights between query and key tokens associated with the object. More precisely, for each latent frame $i$ at diffusion layer $l$, we calculate the attention map as follows:
@@
(3) $A_{i}^{l}=\text{softmax}\left(\frac{Q_{i}^{l}(M_{obj}\odot K_{i}^{l})^{T}}{\sqrt{d}}\right)\in{\mathbb{R}}^{w\times h}$
@@
[…]
[Description]Description
Figure 4. Self-attention maps from (a) LTX Video diffusion model and (b) Stable Diffusion (image based). The spatio-temporal video latent ”attends to object associated effects” (e.g., shadow, reflection) where, image models struggles to capture these associations.
We then compute a mask per-frame, $\mathcal{M}_{i}\in[0,1]^{w\times h\times c}$, which is aggregated across all diffusion layers:
(4) $\mathcal{M}_{i}=\frac{1}{L}\sum_{l=1}^{L}A_{i}^{l}$
where $L$ is the number of layers in the video diffusion model. Then we extend $\mathcal{M}_{i}$ to the whole video by concatenating all frame masks channel-wise $\mathcal{M}_{obj}=\text{concat}\{\mathcal{M}_{1},...\mathcal{M}_{f}\}\in{\mathbb{R}}^{w\times h\times f}$ deriving a soft-mask latent. To obtain a binary mask we perform Otsu thresholding (Otsu, 1979) for each latent frame, $\hat{M}_{obj}=\mathbb{I}(\mathcal{M}_{obj})$. This new mask replaces the mask $M_{obj}$ provided as input by the user. Figure 4(a) shows self-attention maps from LTX-Video (HaCohen et al., 2024) of two video frames. The attention maps reveal the model’s ability to localize not only primary objects—like the cat and the walking person—but also their associated physical effects, such as reflections and shadows. This demonstrates the model’s robustness in consistently tracking these cues even when similar visual patterns are present elsewhere in the scene.
Interestingly, to the best of our knowledge, this approach has not been explored for masking object effects (e.g shadows) in images. Unlike video diffusion models, image models do not capture object effects from still images (Winter et al., 2024). Figure 4(b) illustrates this by showing self-attention maps extracted using StableDiffusion (Rombach et al., 2022), a text-to-image diffusion model, which demonstrates that the object does not attend to its associated effects. We believe that this aligns with the principle of common fate in Gestalt psychology (Köhler, 1992), which suggests that elements moving together are perceived as part of the same object. Video diffusion models seem to implicitly leverage this principle by grouping objects with their effects through motion.
[…]
[Description]Description
Figure 5. (a) Foreground Extraction: The target object is extracted by latent code arithmetic, subtracting the background video encoding from the object+background latent (Latent Diff). This initially results in distortions, which are later corrected by replacing pixel values using the original video and a user-provided mask (Latent Diff + Pixel injection). (b) Layer Composition: The extracted object layer is added to a new background latent (Latent Addition). To improve blending, a few steps of noising-denoising are applied, yielding a more natural integration of the object into the new scene (Latent Addition + Refinement). See video examples in the supplemental material
@@
ObjectDrop (Winter et al., 2024) 28.05 0.124 34.22 0.083 31.14 0.104 - -
Video Repaint [LTXVideo]* (Lugmayr et al., 2022) 20.13 0.252 21.15 0.289 20.64 0.271 0 0.4
Video Repaint [Wan2.1]* (Lugmayr et al., 2022) 21.44 0.244 24.16 0.261 22.8 0.253 0 32
Temporal Enhance [LTXVideo]* (Luo et al., 2025) 21.33 0.248 23.01 0.281 22.8 0.265 0 0.04
Temporal Enhance [Wan2.1]* (Luo et al., 2025) 21.93 0.222 24.26 0.253 23.01 0.237 0 3.2
Lumiere inpainting (Bar-Tal et al., 2024) 26.62 0.148 31.46 0.157 29.04 0.153 - 9
Propainter (Zhou et al., 2023) 27.44 0.114 34.67 0.056 31.06 0.085 - 0.083
DiffuEraser (Zhou et al., 2023) 29.51 0.105 35.19 0.048 32.35 0.077 - 0.8
Ominmatte (Lu et al., 2021) 21.76 0.239 26.81 0.207 24.29 0.223 3 2.5
D2NeRF (Wu et al., 2024) - - 34.99 0.113 - - 3 2.2
LNA (Kasten et al., 2021) 23.10 0.129 - - - - 8.5 0.4
OmnimatteRF (Lin et al., 2023) 33.86 0.017 40.91 0.028 37.38 0.023 6 3.5
Generative Omnimatte (Lee et al., 2024) 32.69 0.030 44.07 0.010 38.38 0.020 - 9
OmnimatteZero [LTXVideo] (Ours) 35.11 0.014 44.97 0.010 40.04 0.012 0 0.04
OmnimatteZero [Wan2.1] (Ours) 34.12 0.017 45.55 0.006 39.84 0.011 0 3.2
Table 1. Quantitative comparison: Background Reconstruction. OmnimatteZero outperforms all omnimatte and video inpainting methods, achieving the best PSNR and LPIPS without training or per-video optimization. It also runs significantly faster, with OmnimatteZero [LTXVideo] at 0.04s per frame. ”-” denotes missing values due to unreported data or unavailable public code.
5.3. Foreground extraction
We can now use our object removal approach to extract object layers along with their associated effects. To isolate a specific object, we apply our method twice: first, removing all objects from the video, leaving only the background, denoted as $V_{bg}\in\mathbb{R}^{F\times W\times H\times 3}$ with its corresponding latent $Z_{bg}$; second, removing all objects except the target object, resulting in a video of the object with the background, denoted as $V_{obj+bg}\in\mathbb{R}^{F\times W\times H\times 3}$ with its corresponding latent $Z_{obj+bg}$. We can now derive the video of the object isolated from the background by simply subtracting the two latents $Z_{obj}=Z_{obj+bg}-Z_{bg}$. Applying thresholding on $Z_{obj}$ results in a latent that is decoded to a video $V_{obj}$ containing only the object and its associated effects (see Figure 5a Latent Diff). While the extracted effects appear convincing, the object itself often suffers from distortions due to latent values falling outside the diffusion model’s learned distribution. To address this issue, we make use of the information available in the pixel-space. We refine the object’s appearance by replacing its values in the pixel-space with those from the original video, based on the user provided mask:
(5) $V_{obj}=M^{p}_{obj}\odot V_{obj+bg}+\left(1-M^{p}_{obj}\right)\odot V_{obj}.$
@@ 5.4. Layer composition
6. Experiments
Following (Lee et al., 2024), we evaluate OmnimatteZero on three applications: (1) Background Reconstruction, where the foreground is removed to recover a clean background; (2) Foreground Extraction, where objects are extracted, together with their associated effects (e.g. shadows and reflections); and (3) Layer Composition, where extracted elements are reinserted into new backgrounds while preserving visual consistency. For qualitative results, each figure shows one representative frame per video; full videos are in the supplementary material.
[…]
[Description]Description
Figure 6. Qualitative Results: Object removal and background reconstruction. The first row shows input video frames with object masks, while the second row presents the reconstructed backgrounds. Our approach effectively removes objects while preserving fine details, reflections, and textures, demonstrating robustness across diverse scenes. Notice the removal of the cat’s reflection in the mirror and water, the shadow of the dog and bicycle (with the rider), and the bending of the trampoline when removing the jumpers. See video examples in the supplemental material
6.1. Background layer reconstruction
@@
We conducted a comparative evaluation of our approach with several SoTA methods. These methods fall into four categories: (A) Omnimatte methods that are trained to decompose a given video into layers: Omnimatte (Lu et al., 2021), D2NeRF (Wu et al., 2024), LNA (Kasten et al., 2021), OmnimatteRF (Lin et al., 2023) and Generative Omnimatte (Lee et al., 2024). (B) Video inpainting methods that are trained to remove objects from a video: RePaint (Lugmayr et al., 2022) adapted for video, Temporal Enhance (Luo et al., 2025) applied to vanilla inpainting (Sec 2) which enhances general temporal attention during denoising, Lumiere inpainting (Bar-Tal et al., 2024), Propainter (Zhou et al., 2023), DiffuEraser (Li et al., 2025), and VideoPainter (Bian et al., 2025). Finally, (C) An image inpainting method that is applied for each video frame independently: ObjectDrop (Winter et al., 2024).
@@
Datasets and Metrics: We evaluate our method on two standard Omnimatte datasets: Movies (Lin et al., 2023) and Kubric (Wu et al., 2024). These datasets provide object masks for each frame and ground truth background layers for evaluation. All methods are assessed using PSNR and LPIPS metrics to measure the accuracy of background reconstruction, along with comparisons of training and runtime efficiency on a single A100 GPU.
@@
Quantitative results: Table 1 presents a comparison of OmnimatteZero with the SoTA Omnimatte and inpainting methods on the Movies and Kubric benchmarks. It shows that OmnimatteZero outperforms existing supervised and self-supervised methods designed for object inpainting or omnimatte. It achieves the highest PSNR and lowest LPIPS on both the Movie and Kubric datasets, demonstrating superior background reconstruction. Specifically, OmnimatteZero [LTXVideo] achieves a PSNR of 35.11 (Movie) and 44.97 (Kubric), surpassing Generative Omnimatte (Lee et al., 2024), all while requiring no training or per-video optimization. Notably, this improvement is not due to a stronger video generation model, as both OmnimatteZero and Video RePaint (Lugmayr et al., 2022) use the same generator, yet Video RePaint records the lowest PSNR and highest LPIPS across all benchmarks.
@@
Our method is also significantly faster. OmnimatteZero [LTXVideo] runs at just 0.04s per frame (or 24 frames per second). These results establish OmnimatteZero as the first training-free, real-time video matting method, offering both superior performance and efficiency.
[…]
[Description]Description
@@
Figure 7. Qualitative Comparison: Foreground Extraction. Foreground extraction comparison between OmnimatteZero and OmnimatteRF (Lin et al., 2023). Our method accurately captures both the object and its associated effects, such as shadows and reflections, in contrast to OmnimatteRF, missing or distorting shadows (row 1) and reflections (row 2).
@@
Qualitative results: Figure 6 presents qualitative results of our object removal method across various scenes. The first row shows input video frames with masked objects, while the second row displays the reconstructed backgrounds. The videos include (1) a cat running toward a mirror, (2) a static cat looking left and right, (3) a dog running toward a man, (4) two people walking, and (5) multiple people jumping on a trampoline. Our method effectively removes objects like people and animals even when similar objects and effects appear in the video, while preserving fine details and textures. Notably, in the first two columns, OmnimatteZero eliminates the cat without leaving mirror or water reflections. The last column further demonstrates its ability to remove objects while maintaining scene integrity, even correcting the trampoline’s bending effect after removing the jumper. Figure 9 10 presents a qualitative comparison of OmnimatteZero with SoTA omnimatte and video inpainting methods. Our method consistently produces cleaner and more temporally coherent background reconstructions with fewer artifacts and distortions, especially in challenging scenes involving motion, shadows, and texture continuity. The figures includes diverse videos such as dogs running on grass, a person walking past a building, a woman walking on the beach, swans in a lake and more. In the first three rows of Figure 10, OmnimatteZero reconstructs textured and complex backgrounds (e.g., grass, water, sand) without the distortions and ghosting present in OmnimatteRF and DiffuEraser (see red boxes). In all rows of Figure 9, competing methods leave visible shadow remnants, semantic drift, or object traces, while OmnimatteZero removes foregrounds cleanly and fills in semantically appropriate content.
[…]
[Description]Description
@@
Figure 8. Qualitative Comparison: Layer Composition. The extracted foreground objects, along with their shadows and reflections, are seamlessly integrated into diverse backgrounds, demonstrating the effectiveness of our approach in preserving visual coherence and realism across different scenes. See video examples in the supplemental material
6.2. Foreground extraction
@@
We aim to extract the foreground object along with its associated effects, such as shadows and reflections. Figure 7 compares OmnimatteZero for foreground extraction with OmnimatteRF (Lin et al., 2023). Our training-free method accurately isolates objects while preserving fine details of their interactions with the scene, such as reflections and shadows. In contrast, OmnimatteRF struggles to fully capture the associated effects. Notably, in the second and third rows, our method correctly extracts both the shadow of the cyclist and the reflection of the cat, while OmnimatteRF either distorts or omits these elements. These results demonstrate that OmnimatteZero provides superior, training-free foreground extraction, effectively handling complex visual effects where existing methods fall short.
6.3. Layer composition